# Three-Torsion Subgroups and Wild Conductor Exponents of Plane Quartics

### v3

@@ Abstract.
In this paper we give an algorithm to find the 3-torsion subgroup of the Jacobian of a smooth plane quartic curve with a marked rational point. We describe $3-$torsion points in terms of cubics which triply intersect the curve, and use this to define a system of equations whose solution set corresponds to the coefficients of these cubics. We compute the points of this zero-dimensional, degree $728$ scheme first by approximation, using homotopy continuation and Newton-Raphson, and then using continued fractions to obtain accurate expressions for these points. We describe how the Galois structure of the field of definition of the $3$-torsion subgroup can be used to compute local wild conductor exponents, including at $p=2$.
@@ 1. Introduction
In this paper we study the $3$-torsion subgroup of genus $3$, non-hyperelliptic curves. We begin by showing that for plane quartics over a number field $K$, which have a $K$-rational point, $3$-torsion points on the Jacobian correspond to cubics of the form
@@
This paper is organised as follows. In Section 2 we give a description of $3$-torsion points using sections of the canonical class and a marked rational point. In Section 3 we describe a method for computing $3$-torsion points. We derive a system of equations whose solutions correspond to $3$-torsion points, which we solve by first approximating the solutions, as tuples of complex numbers, and then by using such approximations and lattice reduction/continued fractions to obtain algebraic expressions for our torsion points. We give an overview of some examples in Section 4. In Section 5 we describe the local Galois representation of $J[3]$ can be computed from our data. In the final section we explain how our calculations can be used to determine wild conductor exponents. We give a brief overview of the description of $2$-torsion subgroup; and explain how this, and consequently, the wild conductor exponents away from $2$, can computed efficiently.
@@
For the rest of this paper, unless stated otherwise, we write $C$ for a smooth and projective, non-hyperelliptic curve of genus $3$, defined over a number field $K$. The curve is embedded in $\mathbb{P}^{2}$ by the canonical embedding and it’s described by a single degree $4$, homogeneous polynomial $f\in K[x,y,z]$. We write $J$ for the Jacobian of $C$. Recall that this is a $3$-dimensional abelian variety over $K$, whose points correspond to linear equivalence classes of degree 00 divisors on the curve, by the Picard functionality.
@@ Acknowledgements.
The authors are grateful to Raymond van Bommel, Tim Dokchitser, Samir Siksek and Damiano Testa for helful conversations. We thank the anonymous referees for their careful reading of this paper. Their comments and suggestions have improved this work greatly. Both authors are grateful for the financial support provided by the UK Engineering and Physical Sciences Research Council. The first named author is supported by the EPSRC Doctoral Prize fellowship EP/W524335/1. The second named author is supported by the EPSRC studentship EP/W523793/1.
@@ 2. Three-Torsion Points and Sections of $K_{C}$
The group of $2$-torsion points is classically generated by differences odd theta characteristics of the curve, which in turn correspond to bitangent lines to the curve, as described in Section 6.3. The problem of describing $3$-torsion points on $J$ is not as straightforward. Notably, a ‘nice geometric description’ analogous to bitangents for $3$-torsion points does not exist, since $\text{GSp}_{6}(\mathbb{F}_{3})$ has no small index subgroups, with the smallest having order $364$, and corresponding to ‘projective‘ $3$-torsion points. In this section we give an elementary description of $3$-torsion points, using sections of the canonical divisor.
@@
Fix an effective representative $K$ of the canonical divisor whose support contains $P_{0}$. For any $D\in J[3]\setminus\{0\}$, we note that $D+K$ is a divisor of degree $4$ and thus by the Riemann-Roch theorem, there exists a unique effective, degree $4$ divisor $D_{0}$, whose support contains $P_{0}$ and such that $D+K_{C}\sim D_{0}$. Thus, there exists non-constant function on the curve $h$, satisfying
$\operatorname{div}h=3D_{0}-3K$.
As $D_{0}$ and $K$ are assumed to be effective, $h\in\mathcal{L}(3K)$ and hence, it is precisely a cubic. As this cubic passes through $P_{0}$, the coefficient of $y^{3}$ is necessarily zero. Moreover, as $z=0$ is the tangent line at this point, we have the additional condition that the coefficient of $y^{2}x$ is also zero. Lastly, the fact that the cubic intersect the quartic with multiplicity $3$ gives us the relation between the coefficient of $x^{2}y$ and $y^{2}z$. ∎
We must note that not all cubics of this form give rise to non-trivial torsion elements. With notation as in the above proof, if $D_{0}-K\sim 0$, then $D_{0}$ is a canonical divisor and so the intersection of $C$ with a line. Since it contains $P_{0}$, the only lines through this point are the lines of constant $x/z$, hence cubics of the form $(x-az)^{3}$ must be ruled out.
Moreover, representations of $3$-torsion points by cubics of the above form are almost always unique. If $D_{0}-P_{0}$ is not a special divisor and $D_{0}-K\sim D_{1}-K$, then $D_{0}-P_{0}\sim D_{1}-P_{0}$, and so there is equality of divisors as neither moves. As the divisors are the same, the cubics cutting them out must be the same up to scaling. If $D_{0}-P_{0}$ is a special divisor, then it can be written as $K^{\prime}-Q$, for some canonical divisor $K^{\prime}$ and point $Q$. It follows that, up to linear equivalence, $D_{0}=P_{0}-Q$, and so $3D_{0}=3P_{0}-3Q\sim 0$. This shows that both $P_{0}$ and $Q$ are flexes. The tangent at $Q$ meets $C$ at another point, $R$, and the divisors equivalent to $3Q$ are the intersection of $C$ with lines through $R$. As $3P_{0}\sim 3Q$, the tangent to $P_{0}$ must meet $C$ again at $R$. So as long as $C$ has either a rational point which is not a flex, or a rational flex whose flex line does not meet another in this way, this method will compute all non-trivial three torsion, and find unique expressions for each element.
@@
We note that this description of torsion can be extended to higher torsion orders and genera: fix a rational degree $g-2$ effective divisor, $\delta$, an effective representative of the canonical, $K$, such that $\delta$ is in the support of $K$, then for any $n$-torsion divisor, $D$ associate to it an effective divisor $D_{0}$ such that $D+K\sim D_{0}$ and $\delta$ is in the support of $D_{0}$. These divisors $D_{0}$ can then be identified with the sections of $nK$ whose vanishing orders are multiples of $n$. For plane degree $d$ curves, this is equivalent to intersecting the curve with degree $n(d-3)$ curves, and for a canonically embedded curve this is intersections with degree $n$ hyperplanes.
@@ 3.1. Schemes of $3$-torsion points
Substituting $y=-s_{3}(x)/s_{2}(x)$ in the equation defining the cubic gives a degree $9$ polynomial $s_{4}(x)$. If the cubic intersects the curve in $3$ points, each occurring with multiplicity $3$, then $s_{4}$ is necessarily a cube, and we set
@@ 3.2. Complex Approximations
Approximate complex solutions of a system as above can be computed by homotopy continuation; a method of obtaining approximate solutions to a system of equations by deforming the known solutions of a similar system. A detailed explanation of this theory can be found in [29]. Homotopy continuation is efficiently implemented in the Julia package HomotopyContinuation.jl (see [3]), and this implementation gives approximates solutions which are accurate to 16 decimal places. We improve the precision of these approximations by the classical method of Newton-Raphson, see [25, Page 298] for a comprehensive overview of this.
@@ 3.3. Lattice Reduction and Precise Expressions
When the corresponding cubic is defined over a field of small degree, we use lattice reduction to determine the minimal polynomials of the coefficients. This is a standard method, see [23, Chapter 6] for details. We give a brief overview of this here.
@@
Assume $\alpha$ is accurate to $k$ decimal places, and fix $k^{\prime}<k$ such that $|[10^{k^{\prime}}\alpha^{i}]-10^{k^{\prime}}x^{i}|\leq 1$, where [][][ ] denotes the closest (Gaussian) integer. If $10^{k^{\prime}}$ is much larger than the coefficients $c_{i}$, then the shortest vector in the lattice spanned by the columns of the following matrix should be $(c_{n},\ldots,c_{1},\gamma)$ where $\gamma=c_{n}[10^{k^{\prime}}\alpha^{n}]+\ldots+c_{1}[10^{k^{\prime}}\alpha]+c_{0}$.
@@
We proceed essentially as before. Let $1,x_{1},\ldots,x_{n}$ be a basis for $K$, and $\alpha$ be a complex approximation to $x$, where $c_{n+1}x=c_{0}+c_{1}x_{1}+\ldots+c_{n}x_{n}$, with $c_{i}\in\mathbb{Z}$, and $\alpha$ is accurate to $k$ decimal places. If $k^{\prime}<k$ is such that $10^{k^{\prime}}$ is much larger than the $c_{i}$, then $(c_{n+1},c_{n},\ldots,c_{1},\gamma)$ will be the shortest vector in the lattice spanned by the columns of the following matrix, where $\gamma=-c_{n+1}[10^{k^{\prime}}\alpha]+c_{n}[10^{k^{\prime}}x_{n}]+\ldots+c_{1}[10^{k^{\prime}}x_{1}]+c_{0}[10^{k^{\prime}}]$.
@@ 3.4. Precise Expressions and Continued Fractions
Let $y_{1},...,y_{d}$ be complex approximations to the roots of a rational monic polynomial of degree $d$. Using Vieta’s formulae [13], we can construct a monic polynomial of degree $d$ whose roots are the $y_{i}$. This polynomial will be an approximation to the original polynomial, and so it suffices to recognise the coefficients as rational numbers.
@@
Unfortunately, it is very hard to check this method produces the correct polynomial as they are typically of extremely large degree. There are some heuristics that can provide evidence, for example, checking whether the number of roots of the polynomial over $\mathbb{F}_{p}$ equals the order of 3-torsion.
@@ 4.1. The $3$-torsion Subgroup of the Fermat Quartic
The transformation $\left(x:y:z\right)\mapsto\left(y:x:y-z\right)$ followed by $\left(x,y,z\right)\mapsto\left(x:\frac{1}{2}y-z:-2z\right)$ moves the rational point $(1:0:1)$ to $P_{0}=(0:1:0)$ with tangent line $z=0$, and we obtain the following model of the curve
@@
We work on the affine patch $z=1$. In this example we were able to use lattice reduction to compute the minimal polynomials of the coefficients of the required cubics. We found $3$ orbits of $3$-torsion points. First, those cut out by cubics of the form
$-y^{2}+uxy+(72u^{6}-14u^{2})y+(1/864)(u^{7}-36u^{3})x^{3}$
where $u^{8}-(1/6)u^{4}-1/432$. The second orbit consists of the cubics of the form
$-y^{2}+uxy+s_{1}(u)y+s_{2}(u)x^{3}+s_{3}(u)x^{2}+s_{4}(u)x+s_{5}(u)$
where $u$ is a solution of $x^{8}-8x^{7}+28x^{6}-56x^{5}+52x^{4}+16x^{3}-80x^{2}+64x-44$ and $s_{1},\ldots,s_{5}$ are explicitly given by
@@ 4.3. Generic Examples
5. Local Galois Representations
@@
For the rest of this section, we assume that $C$ is a plane quartic over $\mathbb{Q}$ and we assume that $C(\mathbb{Q})\neq\emptyset$ as in Section 2. The method described in Section 3 computes a polynomial $F$ of degree $d$, whose points correspond to $3$-torsion points. We view $C$ as a curve over $\mathbb{Q}_{l}$, and note that there is a natural action of $\operatorname{Gal}(\bar{\mathbb{Q}}_{l}/\mathbb{Q}_{l})$ on the roots of $F$, which corresponds to the Galois action on $J[3]$. In this section, we explain how one can recover the Galois representation $\rho:\operatorname{Gal}(\bar{\mathbb{Q}}_{l}/\mathbb{Q}_{l})\rightarrow\operatorname{Aut}(J[3])$ for any $l\neq 2$, from the Galois action on the roots of $F$. We use this local representation to determine the wild conductor exponent at any prime $l\neq 3$, as we shall explain in the next section.
@@ 5.1. Splitting Field and Galois Group
Given $F$, we compute its splitting field $K$ over $\mathbb{Q}_{l}$, the set of roots $R$ over $K$ and the Galois group $G$ of $F$ acting on $R$. Due to the nature of these polynomials, they are often highly reducible and can therefore be factorised more efficiently by naive means, rather than Magma’s implemented local fields package. Our implementation of this is given in the repository, where we compute these factorisations via Hensel lifting and Newton polygons. Finding all the roots of $F$ using Hensel lifting can be intensive, as many start solutions are needed over highly ramified extensions. We exploit the factorisable structure to compute one for each irreducible factor, and then compute its Galois orbit.
Finally, we must compute the permutation action of the Galois group on the roots. The naive approach of computing this requires computing the differences between the image of each root under Galois, and all the other roots. Arithmetic in large field extensions is slow, so we use their images in the residue field to quickly perform comparisons.

### v4

@@ Abstract.
In this paper we give an algorithm to find the 3-torsion subgroup of the Jacobian of a smooth plane quartic curve with a marked rational point. We describe 33-torsion points in terms of cubics which triply intersect the curve, and use this to define a system of equations whose solution set corresponds to the coefficients of these cubics. We compute the points of this zero-dimensional, degree $728$ scheme first by approximation, using homotopy continuation and Newton-Raphson, and then using continued fractions to obtain accurate expressions for these points. We describe how the Galois structure of the field of definition of the $3$-torsion subgroup can be used to compute local wild conductor exponents, including at $p=2$.
@@ 1. Introduction
In this paper we study the $3$-torsion subgroup of genus $3$, non-hyperelliptic curves. We begin by showing that for plane quartics over a number field $K$, which have a $K$-rational point, after a suitable normalisation, $3$-torsion points on the Jacobian correspond to cubics of the form
@@
This paper is organised as follows. In Section 2 we give a description of $3$-torsion points using sections of the canonical class and a marked rational point. In Section 3 we describe a method for computing $3$-torsion points. We derive a system of equations whose solutions correspond to $3$-torsion points, which we solve by first approximating the solutions, as tuples of complex numbers, and then by using such approximations and lattice reduction/continued fractions to obtain algebraic expressions for our torsion points. We give an overview of some examples in Section 4. In Section 5 we describe how the local Galois representation of $J[3]$ can be computed from our data. In Section6 we give a brief overview of the description of the $2$-torsion subgroup, and explain how one compute the local Galois representation at $2$. In the final section we explain how our calculations can be used to determine the wild conductor exponents at all primes.
@@
For the rest of this paper, unless stated otherwise, we write $C$ for a smooth and projective, non-hyperelliptic curve of genus $3$, defined over a number field $K$. The curve is embedded in $\mathbb{P}^{2}$ by the canonical embedding and it’s described by a single degree $4$, homogeneous polynomial $f\in K[x,y,z]$. We write $J$ for the Jacobian of $C$. Recall that this is a $3$-dimensional abelian variety over $K$, whose points correspond to linear equivalence classes of degree 0 divisors on the curve, by the Picard functionality.
@@ Acknowledgements.
The authors are grateful to Raymond van Bommel, Tim Dokchitser, Samir Siksek and Damiano Testa for helpful conversations. We thank the anonymous referees for their careful reading of this paper. Their comments and suggestions have improved this work greatly. Both authors are grateful for the financial support provided by the UK Engineering and Physical Sciences Research Council. The first named author is supported by the EPSRC Doctoral Prize fellowship EP/W524335/1. The second named author is supported by the EPSRC studentship EP/W523793/1.
@@ 2. Three-Torsion Points and Sections of $K_{C}$
The group of $2$-torsion points is classically generated by differences of odd theta characteristics of the curve, which in turn correspond to bitangent lines to the curve, as described in Section 6. The problem of describing $3$-torsion points on $J$ is not as straightforward. Notably, a ‘nice geometric description’ analogous to bitangents for $3$-torsion points does not exist, since $\text{GSp}_{6}(\mathbb{F}_{3})$ has no small index subgroups, with the smallest having order $364$, and corresponding to ‘projective‘ $3$-torsion points. In this section we give an elementary description of $3$-torsion points, using sections of the canonical divisor.
@@
Fix an effective representative $K_{C}$ of the canonical divisor whose support contains $P_{0}$. For any $D\in J[3]\setminus\{0\}$, we note that $D+K_{C}$ is a divisor of degree $4$ and thus by the Riemann-Roch theorem, there exists an effective, degree $4$ divisor $D_{0}$, whose support contains $P_{0}$ and such that $D+K_{C}\sim D_{0}$. Thus, there exists a non-constant function on the curve $h$, satisfying
$\operatorname{div}h=3D_{0}-3K_{C}$.
As $D_{0}$ and $K_{C}$ are assumed to be effective, $h\in\mathcal{L}(3K_{C})$ and hence, it is precisely a cubic. As this cubic passes through $P_{0}$, the coefficient of $y^{3}$ is necessarily zero. Moreover, as $z=0$ is the tangent line at this point, we have the additional condition that the coefficient of $y^{2}x$ is also zero. Lastly, the fact that the cubic intersect the quartic with multiplicity $3$ gives us the relation between the coefficient of $x^{2}y$ and $y^{2}z$. Working in the affine patch obtained by dehomogenising with respect to $y$ shows that at $P_{0}$, $x$ is a uniformiser and $z=\alpha_{f}x^{2}\mod x^{3}$, and for the cubic to vanish to order 3 at $P_{0}$, this equality must hold there too. ∎
We must note that not all cubics of this form give rise to non-trivial torsion elements. With notation as in the above proof, if $D_{0}-K_{C}\sim 0$, then $D_{0}$ is a canonical divisor and so the intersection of $C$ with a line. Since it contains $P_{0}$, the only lines through this point are the lines of constant $x/z$, hence cubics of the form $(x-az)^{3}$ are excluded from the above correspondence.
Moreover, representations of $3$-torsion points by cubics of the above form are almost always unique. If $D_{0}-P_{0}$ is not a special divisor (that is, $l(K_{C}-(D_{0}-P_{0}))=0$) and $D_{0}-K_{C}\sim D_{1}-K_{C}$, then $D_{0}-P_{0}\sim D_{1}-P_{0}$, and so there is equality of divisors as neither moves. As the divisors are the same, the cubics cutting them out must be the same up to scaling. If $D_{0}-P_{0}$ is a special divisor, then it can be written as $K^{\prime}-Q$, for some canonical divisor $K^{\prime}$ and point $Q$. It follows that, up to linear equivalence, $D\sim P_{0}-Q$, and so $3D\sim 3P_{0}-3Q\sim 0$. This shows that both $P_{0}$ and $Q$ are flexes. The tangent at $Q$ meets $C$ at another point, $R$, and the divisors equivalent to $3Q$ are the intersection of $C$ with lines through $R$. As $3P_{0}\sim 3Q$, the tangent to $P_{0}$ must meet $C$ again at $R$. So as long as $C$ has either a rational point which is not a flex, or a rational flex whose flex line does not meet another in this way, this method will compute all non-trivial three torsion, and find unique expressions for each element.
@@
We note that this description of torsion can be extended to higher torsion orders and genera: fix a rational degree g2g-2 effective divisor, $\delta$, an effective representative of the canonical, $K$, such that $\delta$ is in the support of $K$, then for any $n$-torsion divisor, $D$ associate to it an effective divisor $D_{0}$ such that $D+K\sim D_{0}$ and $\delta$ is in the support of $D_{0}$. These divisors $D_{0}$ can then be identified with the sections of $nK$ whose vanishing orders are multiples of $n$. For plane degree $d$ curves, this is equivalent to intersecting the curve with degree n(d3)n(d-3) curves, and for a canonically embedded curve this is intersections with degree $n$ hyperplanes.
@@ 3.1. Schemes of $3$-torsion points
Substituting $y=-s_{3}(x)/s_{2}(x)$ in the equation defining the cubic gives a rational function whose numerator is a degree $9$ polynomial $s_{4}(x)$. If the cubic intersects the curve in $3$ points, each occurring with multiplicity $3$, then $s_{4}$ is necessarily a cube, and we set
@@ 3.2. Complex Approximations
Approximate complex solutions of a system as above can be computed by homotopy continuation; a method of obtaining approximate solutions to a system of equations by deforming the known solutions of a similar system. A detailed explanation of this theory can be found in [31]. Homotopy continuation is efficiently implemented in the Julia package HomotopyContinuation.jl (see [3]), and this implementation gives approximates solutions which are accurate to 16 decimal places. We improve the precision of these approximations by the classical method of Newton-Raphson, see [27, Page 298] for a comprehensive overview of this.
@@ 3.3. Lattice Reduction and Precise Expressions
When the corresponding cubic is defined over a field of small degree, we use lattice reduction to determine the minimal polynomials of the coefficients. This is a standard method, see [25, Chapter 6] for details. We give a brief overview of this here.
@@
Assume $\alpha$ is accurate to $k$ decimal places, and fix $k^{\prime}<k$ such that $|[10^{k^{\prime}}\alpha^{i}]-10^{k^{\prime}}x^{i}|\leq 1$, where $[]$ denotes the closest (Gaussian) integer. If $10^{k^{\prime}}$ is much larger than the coefficients $c_{i}$, then the shortest vector in the lattice spanned by the columns of the following matrix should be $(c_{n},\ldots,c_{1},\mathrm{Re}(\gamma),\mathrm{Im}(\gamma))$ where $\gamma=c_{n}[10^{k^{\prime}}\alpha^{n}]+\ldots+c_{1}[10^{k^{\prime}}\alpha]+c_{0}10^{k^{\prime}}$.
@@
We proceed essentially as before. Let $1,x_{1},\ldots,x_{n}$ be a basis for $K$, and $\alpha$ be a complex approximation to $x$, where $c_{n+1}x=c_{0}+c_{1}x_{1}+\ldots+c_{n}x_{n}$, with $c_{i}\in\mathbb{Z}$, and $\alpha$ is accurate to $k$ decimal places. If $k^{\prime}<k$ is such that $10^{k^{\prime}}$ is much larger than the $c_{i}$, then $(-c_{n+1},c_{n},\ldots,c_{1},\mathrm{Re}(\gamma),\mathrm{Im}(\gamma))$ will be the shortest vector in the lattice spanned by the columns of the following matrix, where $\gamma=-c_{n+1}[10^{k^{\prime}}\alpha]+c_{n}[10^{k^{\prime}}x_{n}]+\ldots+c_{1}[10^{k^{\prime}}x_{1}]+c_{0}[10^{k^{\prime}}]$.
@@ 3.4. Precise Expressions and Continued Fractions
Let $y_{1},...,y_{d}$ be complex approximations to the roots of a rational monic polynomial of degree $d$. We can then compute the polynomial $\prod_{i}(x-y_{i})$. This polynomial will be an approximation to the original polynomial, and so it suffices to recognise the coefficients as rational numbers.
@@
Unfortunately, it is very hard to check this method produces the correct polynomial as they are typically of extremely large degree. There are some heuristics that can provide evidence, for example, checking whether the number of roots of the polynomial over $\mathbb{F}_{p}$ equals the order of 3-torsion over $\mathbb{F}_{p}$.
@@ 4.1. The $3$-torsion Subgroup of the Fermat Quartic
The transformation $\left(x,y,z\right)\mapsto\left(y,x,y-z\right)$ followed by $\left(x,y,z\right)\mapsto\left(x,\frac{1}{2}y-z,-2z\right)$ moves the rational point $(1:0:1)$ to $P_{0}=(0:1:0)$ with tangent line $z=0$, and we obtain the following model of the curve
@@
We work on the affine patch $z=1$. In this example we were able to use lattice reduction to compute the minimal polynomials of the coefficients of the required cubics. We found $3$ Galois orbits of $3$-torsion points. First, those cut out by cubics of the form
$y^{2}+uxy+(72u^{6}-14u^{2})y+(1/864)(u^{7}-36u^{3})x^{3}$
where $u^{8}-(1/6)u^{4}-1/432=0$. The second orbit consists of the cubics of the form
$y^{2}+uxy+s_{1}(u)y+s_{2}(u)x^{3}+s_{3}(u)x^{2}+s_{4}(u)x+s_{5}(u)$
where $u$ is a solution of $x^{8}-8x^{7}+28x^{6}-56x^{5}+52x^{4}+16x^{3}-80x^{2}+64x-44=0$ and $s_{1},\ldots,s_{5}$ are explicitly given by
@@ 4.3. Generic Examples
5. Local Galois Representations at 33
@@
For the rest of this section, we assume that $C$ is a plane quartic over $\mathbb{Q}$ and we assume that $C(\mathbb{Q})\neq\emptyset$ as in Section 2. The method described in Section 3 computes a polynomial $F$ of some degree $d$, whose points correspond to $3$-torsion points. We view $C$ as a curve over $\mathbb{Q}_{\ell}$, and note that there is a natural action of $\operatorname{Gal}(\bar{\mathbb{Q}}_{\ell}/\mathbb{Q}_{\ell})$ on the roots of $F$, which corresponds to the Galois action on $J[3]$. In this section, we explain how one can recover the Galois representation $\rho:\operatorname{Gal}(\bar{\mathbb{Q}}_{\ell}/\mathbb{Q}_{\ell})\rightarrow\operatorname{Aut}(J[3])$ for any $\ell\neq 2$, from the Galois action on the roots of $F$. We use this local representation to determine the wild conductor exponent away from $3$.
@@ 5.1. Splitting Field and Galois Group
Given $F$, we compute its splitting field $K$ over $\mathbb{Q}_{\ell}$, the set of roots $R$ over $K$ and the Galois group $G$ of $F$ acting on $R$. In Magma, this is done using the algorithms given in [13] and [23]. In some of our examples, we found it more efficient to compute the splitting field directly. Our implementation of this is given in the repository, where we compute these factorisations via Hensel lifting and Newton polygons. Finding all the roots of $F$ using Hensel lifting can be intensive, as many start solutions are needed over highly ramified extensions. We exploit the factorisable structure to compute one for each irreducible factor, and then compute its Galois orbit.
The Galois group is computed from the splitting field. Finally, we must compute the permutation action of the Galois group on the roots. The naive approach of computing this requires computing the differences between the image of each root under Galois, and all the other roots. Arithmetic in large field extensions is slow, so we use their images in the residue field to quickly perform comparisons.